Find a number such that four times two more than the number is the same as seven less than the opposite of the number.
Answer by mananth(16949)

let the number be x
4 times the number = 4x
two more = 4x+2
..
opposite number = -x
7 less = -x-(-7)
..
4x+2=-x+7
4x+x=7-2
5x=5
x=1